# AutoHFormer: Efficient Hierarchical Autoregressive Transformer for Time Series Prediction

### v1

@@ V Conclusion
We present AutoHFormer, a hierarchical autoregressive transformer that fundamentally advances time series forecasting by simultaneously addressing three critical challenges: strict causality, computational efficiency, and multi-scale pattern recognition. The architecture integrates three key innovations: (1) A hierarchical processing framework combining segment-level parallel computation (for long-range dependencies) with stepwise sequential refinement (for local precision), (2) Dynamic windowed attention employing learnable exponential decay and adaptive window sizing to maintain temporal coherence while reducing computational overhead, and (3) Hybrid temporal encodings blending sinusoidal patterns (for short-term variations) with learned decay rates (for long-term trends). Extensive experiments demonstrate AutoHFormer’s superiority over PatchTST, achieving 10.76× faster training speeds and 6.06× memory reduction on PEMS8 benchmarks while preserving accuracy across 96–720-step horizons. The system’s strict causal constraints, parallelizable design, and multi-scale modeling capabilities make it particularly suited for industrial applications requiring both efficiency and precision.

### v2

@@ V Conclusion
We present AutoHFormer, a hierarchical autoregressive transformer that advances time series forecasting by addressing three key challenges: strict causality, computational efficiency, and multi-scale pattern recognition. The architecture introduces a hierarchical framework combining segment-level parallel computation for long-range dependencies with stepwise refinement for local precision. Experiments show AutoHFormer’s superiority over PatchTST.
VI Acknowledgements
This research is partly supported by HKU-SCF FinTech Academy, Shenzhen-Hong Kong-Macao Science and Technology Plan Project (Category C Project: SGDX20210823103537030), Theme-based Research Scheme of RGC, Hong Kong (T35-710/20-R).
AI-generated content disclosure acknowledgement
This paper utilized generative AI tools exclusively to improve the linguistic clarity and quality of the manuscript, in a manner comparable to conventional writing aids such as Grammarly. All substantive intellectual contributions, including theoretical formulation, methodological design, experimental implementation, data analysis, and interpretation of findings, are entirely the work of the authors. The AI tools served solely as advanced proofreading assistants and did not contribute to any conceptual or analytical aspects of the research.